Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)
A. y = x
B. y = x + 6
C. y = x – 6
D. y = x + 6

Answer by jim_thompson5910(35256)

y = x + 8 ----> m = 1, b = 8

-------------------------------------------------------

y = mx+b

y = 1x + b

y = x + b

10 = 4 + b

10-4 = b

6 = b

b = 6

-------------------------------------------------------

y = mx+b

y = 1x + 6

y = x + 6

So the answer is choice B